Kay can eat 32 candies in 1 hour. Her sister can eat the same amount in 3 hours. How long will it take them together to eat the 32 pieces of candy.
Answer by ptaylor(1894)


Let x=amount of time it takes for Kay and Sis to eat 32 pieces of candy
Now we know that:
Kay eats candies at the rate of 32 candies per hour.
Sis eats candies at the rate of 32/3 candies per hour.
Kay and Sis together can eat: 32+32/3 or (96+32)/3=128/3 candies per hour
Now we can set up a ratio:
128/3 candies is to 1 hour as 32 candies is to x hours and we can write this as follows:
(128/3)/1=32/x or
128/3=32/x multiply both sides by 3x
128x=96 divide both sides by 128
x=0.75 hr----------------------------------answer
Possibly an easier way to look at this problem is as follows:
If Kay and Sis can eat 128/3 candies per hour, then how many hours will it take them to eat 32 candies? Here's the equation we need to solve:
(128/3)x=32 multiply both sides by 3
128x=96
x=0.75 hr


Hope this helps----ptaylor